# K-Act2Emo: Korean Commonsense Knowledge Graph for Indirect Emotional Expression

### 21 Mar 2024 v1

@@ 7 Discussion
The Level of Volition in Actions The link between emotions and actions, especially in if-then reasoning, is complex. For instance, while it might make sense to say ’PersonX laughs loudly because he feels angry,’ deducing that the laughter is due to anger defies typical logic. This highlights how voluntary actions can skew the true representation of emotions. Our study sought to keep emotional inferences within common-sense bounds, opting for "as if he/she feels" to prevent misassociating actions with emotions, like laughter with anger. This highlights the necessity for finer distinctions in subsequent analyses.
@@ Appendix A Materials for Collection of Indirect Emotional Expressions
The list of emotion-related words from Hong and Jeong (2009) is available on our repository 1 in CSV format.
@@ Appendix B Training Details
Details about COMET-BART Training For COMET-BART training, we utilized the Hugging Face version (hyunwoongko/kobart) of KoBART as a base model. We fine-tuned BART for six epochs using a batch size of 32 and a learning rate of 3e-4 on an NVIDIA A100 GPU.
Details about LLM Generation For both 5-shot and 1-shot learning, we utilized the following list of Korean LLMs and GPT-4 Turbo (0125-preview). We set the temperature and top-p sampling rate to 1.0 each. For the GPT-4 Turbo model, we followed the default hyperparameters suggested by the OpenAI Platform, except for the temperature, which we set to 0.7.
@@
Mi:dm is a 5.7B-sized, instruction-tuned auto-regressive language model. It is trained on both Korean and English in-house dataset. We utilized the Hugging Face version (KT-AI/midm-bitext-S-7B-inst-v1).
ChatSKKU is a 5.8B-sized, instruction-tuned model based on polyglot-ko-5.8b (Ko et al., 2023) (a Korean Large Language Model trained with 172 billion tokens using the GPT-NeoX framework). It is instruction-tuned with the SKKU dataset, which includes a daily Korean conversation dataset. We utilized the Hugging Face version (jojo0217/ChatSKKU5.8B).
KoAlpaca We used a 5.8B-sized version of KoAlpaca based on polyglot-ko-5.8. The KoAlpaca Dataset v1.1b, consisting of a translated version of the Stanford Alpaca dataset and ChatGPT-generated data based on the Naver Q&A dataset, is used for instruction-tuning. We utilized the Hugging Face version (beomi/KoAlpaca-Polyglot-5.8B).
KoPlatypus We used the 7B-sized COT version of KoPlatypus, based on the LLaMA 2 (Touvron et al., 2023) transformer architecture. The KoCoT_2000 dataset, a translated version of the KAIST-CoT dataset, was used for instruction-tuning. We utilized the Hugging Face version (kyujinpy/KoT-platypus2-7B).
Kullm We used a 5.8B-sized version of Kullm based on polyglot-ko-5.8b. The Kullm v2 dataset, which includes translated versions of the GPT-4 LLM, Vicuna, and Dolly datasets, was used for instruction-tuning. We utilized the Hugging Face version (nlpai-lab/kullm-polyglot-5.8b-v2).
Appendix C Fine-Grained Statistics on Emotion Categories of Data Split

### 23 Mar 2024 v2

@@ 6 Discussion
The Level of Volition in Actions The link between emotions and actions, especially in if-then reasoning, is complex. For instance, although it might make sense to say, "PersonX laughs loudly because he feels angry," deducing that the laughter is due to anger defies typical logic. This highlights how voluntary actions can skew the true representation of emotions. Our study sought to keep emotional inferences within common-sense bounds, opting for "as if he/she feels" to prevent misassociating actions with emotions, like laughter with anger. This highlights the necessity for finer distinctions in subsequent analyses.
@@ Appendix A Materials for Collection of Indirect Emotional Expressions
The list of emotion-related words from Hong and Jeong (2009) is available on our repository 1.
@@
Table 15: List of body parts used for collecting indirect emotional expressions.
@@ Appendix B Training Details
Details about COMET-BART Training For COMET-BART training, we utilize the Hugging Face version (hyunwoongko/kobart) of KoBART as a base model. We fine-tune BART for six epochs using a batch size of 32 and a learning rate of 3e-4 on an NVIDIA A100 GPU 40 GB.
Details about LLM Generation For both five-shot and one-shot learning, we utilize the following list of Korean LLMs and GPT-4 Turbo (0125-preview). We set the temperature and top-p sampling rate to 1.0 each. For the GPT-4 Turbo model, we follow the default hyperparameters suggested by the OpenAI Platform, except for the temperature, which we set to 0.7.
@@
Mi:dm is a 5.7B-sized, instruction-tuned auto-regressive language model. It is trained on both Korean and English in-house dataset. We utilize the Hugging Face version (KT-AI/midm-bitext-S-7B-inst-v1).
ChatSKKU is a 5.8B-sized, instruction-tuned model based on polyglot-ko-5.8b (Ko et al., 2023) (a Korean Large Language Model trained with 172 billion tokens using the GPT-NeoX framework). It is instruction-tuned with the SKKU dataset, which includes a daily Korean conversation dataset. We utilize the Hugging Face version (jojo0217/ChatSKKU5.8B).
KoAlpaca We use a 5.8B-sized version of KoAlpaca based on polyglot-ko-5.8. The KoAlpaca Dataset v1.1b, consisting of a translated version of the Stanford Alpaca dataset and ChatGPT-generated data based on the Naver Q&A dataset, is used for instruction-tuning. We utilize the Hugging Face version (beomi/KoAlpaca-Polyglot-5.8B).
KoPlatypus We use the 7B-sized COT version of KoPlatypus, based on the LLaMA 2 (Touvron et al., 2023) transformer architecture. The KoCoT_2000 dataset, a translated version of the KAIST-CoT dataset, is used for instruction-tuning. We utilize the Hugging Face version (kyujinpy/KoT-platypus2-7B).
Kullm We use a 5.8B-sized version of Kullm based on polyglot-ko-5.8b. The Kullm v2 dataset, which includes translated versions of the GPT-4 LLM, Vicuna, and Dolly datasets, is used for instruction-tuning. We utilize the Hugging Face version (nlpai-lab/kullm-polyglot-5.8b-v2).
Templates used in few-shot learning
Table 16 is a basic version of the templates we use for one-shot and five-shot learning. "Positive" can change to "Negative" in the case of NegEnv, and the example is randomly sampled from the training set each time. We adapt the template to the usage guideline of each model. For example, when we use Mi:dm, we replace “Response:” with “Midm;”.
Template
아래는 작업을 설명하는 명령어와 추가적 맥락을 제공하는 입력이
짝을 이루는 예제입니다.
요청을 적절히 완료하는 응답을 작성하세요.
###명령어: 긍정적인 상황 속에서 이 사람의 감정을 추론하라.
한 문장으로 답하라.
###입력 1: Y에게 가까이 다가간다.
###응답 1: 친밀감을 느끼다.
###입력 2: 고개를 갸웃거린다.
###응답 2:
Below is an instruction that describes a task, paired with an input that
provides further context.
Write a response that appropriately completes the request.
###Instruction: Infer this person’s emotion in a positive situation.
Answer in one sentence.
###Input 1: He/she approaches Y closely.
###Response 1: He/she feels intimacy.
###Input 2: He/she tilts his/her head.
###Response 2:
Table 16: Templates used in one-shot learning. For five shot learning, we add four more examples.
Appendix C Statistics on Emotion Categories of Data Split